A billboard is 7 meters long and 5 meters high. An advertiser wishes to increase its size by adding the same amount to its length and width. The new area will be 63 square meters. How much must be added to the length and width of the billboard to do this?
Found 2 solutions by piglet162431, JWG:
Answer by piglet162431(90)

If a original billboard is 7 X 5 meters, then the area will be 35. 63-35= 28. 28 of the length and width must be aded.

Answer by JWG(21)

I'm not sure the above is correct:
Area of a rectangle is length*width, and each side is going to increased by the same amount to get the new area of 63.
(7+x)(5+x)=63
35+12x+x^2=63
x^2+12x+35-63=0
x^2+12x-28=0
(x-2)(x+14)=0
x=2 and x=-14
Only logical solution is x=2
Check: (7+2)(5+2)=9*7. 9*7=63